# MADNESS Deblender

### v1

@@ Appendix A SNR computation
Using the scarlet release btk-v1, all the galaxies were modeled as per standard recommendations in scarlet documentation [^13] https://pmelchior.github.io/scarlet/. The sources are initialized using the scarlet.initialization.init_all_sources() with all parameters set to default except for the max_components which is set to $2$ so that the buldge and disk components can be properly modeled provided that the SNR is large enough. The min_snr parameter parameter corresponding to this SNR value is set to the recommended default value of $50$.
@@ Appendix C Metrics
The blendedness ($\beta$) metric defined in Section 4.1 lies between 0 and 1, 0 for isolated galaxies, and close to 1 for severely blended ones. Figure 16 gives a visual representation of the variation of blendedness for different values of SNR in the r-band.
@@
The contamination ($\kappa$) defined in 4.1 also quantifies the effect of blending. For isolated galaxies, the value of $\kappa$ is 00 and keeps increasing according to the contribution to the flux from neighboring galaxies within the flux radius of aperture-photometry. Figure 17 gives a visual representation of how $\kappa$ varies in the r-band for blended galaxies at different SNRs.
@@
Figure 16: Normalized images of galaxies in r-band for different levels of blendedness $\beta_{r}$ of the central galaxy (centers marked with black $\times$), in the presence of neighboring galaxies (centers marked with red $\times$). From left to right, the value of blendedness increases, and from top to bottom the SNR of the central galaxy increases. The color bar for each plot shows the normalized electron counts in the pixels.
@@
Figure 17: Normalized images of galaxies in r-band for different levels of contamination $\kappa_{r}$ of the central galaxy (centers marked with black $\times$), in the presence of neighboring galaxies (centers marked with red $\times$). From left to right, the value of contamination increases, and from top to bottom, the SNR of the central galaxy increases. The color bar for each plot shows the normalized electron counts in the pixels.
Appendix D Photometry across bands
(a) Photometry results in each LSST band, using all the galaxies in TestFields.
(b) Photometry results in each LSST band, neglecting low SNR galaxies in TestFields.
@@ Appendix D Photometry across bands
Figure 18: Cumulative distribution of the absolute value of the relative flux residuals $|\delta F/F|$ before deblending and after deblending with MADNESS, scarlet, and VAE-deblender, for each LSST band. The results are presented for two cases: (a) all galaxies in TestFields and (b) galaxies with $\text{SNR}_{\text{blend}}\geq 5$.
@@
In Section 4.2 we showed that MADNESS outperformed the flux reconstructions of scarlet in the r-band. In Figure 18 compare the flux reconstructions among MADNESS, scarlet, and VAE-deblender for all the bands of LSST.
@@
Figure 18a shows the results for all the galaxies in the TestFields dataset and we observe that the performance of MADNESS and VAE-deblender drops significantly u-band, compared to that of scarlet. To see if this is due to the low SNR of galaxies in the u-band, we redo the plot with only high SNR galaxies in each of the bands. To simplify the SNR computation in each band, we can define SNRblend for each band in the blended configuration as:
$\text{SNR}_{\text{blend}}=\text{flux}/\text{fluxerr};$ (27)
where flux and flux are obtained from aperture-photometry with the SourceExtractor library. For each band, after discarding galaxies with SNRblend5absent5$\leq$ 5≤ 5 see in Figure 18b that now the performance of MADNESS is comparable to scarlet even in the u-band. Therefore, we can conclude that the low SNR in u-band leads to poor flux reconstructions.
Appendix E SSIM across bands
(a) SSIM results using all the galaxies in TestFields.
(b) SSIM results neglecting low SNR galaxies in TestFields.
Figure 19: Comparison of the structural similarity index measure (SSIM) after deblending with MADNESS, scarlet, and VAE-deblender, for each LSST band. The results are presented for two cases: (a) all galaxies in TestFields and (b) galaxies with $\text{SNR}_{\text{blend}}\geq 5$. A value closer to 00 for $1-\rm{SSIM}$ on the x-axis, represents better performance.
Since we observed in Appendix D that there is a slight dip in the performance of MADNESS flux reconstructions in the u-band, in this Section we want to look into the morphology of the deblended models in different bands. Using SSIM as a metric, we show in Figure 19 that scarlet reconstructs morphology far better than MADNESS in the u-band although in all the other bands MADNESS consistently outperforms scarlet.
@@ Appendix E SSIM across bands
We can try several ways to mitigate this problem. First, we can try to add a weighting function to Equation 15 to enforce better reconstructions in the u-band. As an alternative, we can also adopt a non-linear data normalization like Arcelin et al. (2020) where authors suppress the dominance of brighter pixels using the normalization to increase the relative weights of the fainter pixels. Finally, we can also try to adopt a scarlet-like approach to hierarchically predict the morphology and the SED of galaxies but it would constrain us to the unrealistic assumption of the same morphology in all the bands.

### v2

@@ Appendix C Comparison between u- and r-band
Using the scarlet release btk-v1, all the galaxies were modeled as per standard recommendations in scarlet documentation [^14] https://pmelchior.github.io/scarlet/. The sources are initialized using the scarlet.initialization.init_all_sources() with all parameters set to default except for the max_components, which is set to $2$ so that the bulge and disk components can be properly modeled, provided that the S/N is large enough. The min_snr parameter corresponding to this S/N value is set to the recommended default value of $50$.
@@ Appendix E Metrics
The blendedness ($\beta$) metric defined in Section 4.1 lies between 0 and 1, 0 for isolated galaxies, and close to 1 for severely blended ones. Figure 17 gives a visual representation of the variation of blendedness for different values of S/N in the r-band.
@@
The contamination ($\kappa$) defined in 4.1 also quantifies the effect of blending. For isolated galaxies, the value of $\kappa$ is 0 and increases according to the contribution to the flux from neighboring galaxies within the flux radius of aperture photometry. Figure 18 gives a visual representation of how $\kappa$ varies in the r-band for blended galaxies at different S/Ns.
@@
Figure 17: Normalized images of galaxies in r-band for different levels of blendedness $\beta_{r}$ of the central galaxy (centers marked with black $\times$), in the presence of neighboring galaxies (centers marked with red $\times$). From left to right, the value of blendedness increases, and from top to bottom, the S/N of the central galaxy increases. The color bar for each plot shows the normalized electron counts in the pixels.
@@
Figure 18: Normalized images of galaxies in r-band for different levels of contamination $\kappa_{r}$ of the central galaxy (centers marked with black $\times$), in the presence of neighboring galaxies (centers marked with red $\times$). From left to right, the value of contamination increases, and from top to bottom, the S/N of the central galaxy increases. The color bar for each plot shows the normalized electron counts in the pixels.
Appendix F Photometry across bands
@@
Figure 19: Cumulative distribution of the absolute value of the relative flux residuals $|\delta F/F|$ before deblending and after deblending with MADNESS, scarlet, and VAE-deblender, for each LSST band. The results are presented for two cases: (a) all galaxies in TestFields and (b) galaxies in TestFields with $\text{S/N}_{\text{blend}}\geq 5$.
@@
In Section 4.2, we showed that MADNESS outperformed the flux reconstructions of scarlet in the r-band. In Figure 19, compare the flux reconstructions among MADNESS, scarlet, and VAE-deblender for all the bands of LSST.
@@
Figure 19a shows the results for all the galaxies in the TestFields dataset and we observe that the performance of MADNESS and VAE-deblender drops significantly u-band, compared to that of scarlet. To see if this is due to the low S/N of galaxies in the u-band, we redo the plot with only high S/N galaxies in each of the bands. To simplify the S/N computation in each band, we can define S/Nblend for each band in the blended configuration as:
$\text{S/N}_{\text{blend}}=\texttt{flux}/\texttt{fluxerr};$ (22)
@@
Appendix G SSIM across bands
Figure 20: Comparison of the structural similarity index measure (SSIM) after deblending with MADNESS, scarlet, and VAE-deblender, for each LSST band. The results are presented for two cases: (a) all galaxies in TestFields and (b) galaxies in TestFields with $\text{S/N}_{\text{blend}}\geq 5$. A value closer to 0 for $1-\rm{SSIM}$ on the x-axis represents better performance.
Since we observed in Appendix F that there is a slight dip in the performance of MADNESS flux reconstructions in the u-band, in this Section we want to look into the morphology of the deblended models in different bands. Using SSIM as a metric, we show in Figure 20 that scarlet reconstructs morphology far better than MADNESS in the u-band, although in all the other bands MADNESS consistently outperforms scarlet.
[…]
We can try several ways to mitigate this problem. First, we can try to add a weighting function to Equation 10 to enforce better reconstructions in the u-band. As an alternative, we can also adopt a non-linear data normalization similar to that of Arcelin et al. (2020) whose authors suppressed the dominance of brighter pixels using the normalization to increase the relative weights of the fainter pixels. Finally, we can also try to adopt a scarlet-like approach to hierarchically predict the morphology and the SED of galaxies but it would constrain one to the unrealistic assumption of the same morphology in all the bands.